A 1 kg steamActivity 2 0 water mixture at 1.0 MPa is contained in an inflexible tank. Heat is added until the pressure rises to 3.5 MPa and 400 C. Determine the heat added
1
Expert's answer
2021-08-04T02:52:28-0400

"m = 1 kg\\
P_1 = 1.0\ MPa\\
P_2 = 3.5\ MPa\\
U_1 = 2401.41\frac{ kJ} {kg}\\
U_2= 2926.4 \frac{ kJ} {kg}\\
T = 400°C"
Heat transferred = "m(U_2-U_1) =""1(2926.4-2401.41) = 524.99""= 524.99 kJ"